# New Sorting Algorithm Wave Sort (W-Sort)

### v1

@@ 2. Core Algorithm
Once the down-wave completes sorting this set, thereby enlarging the overall sorted region, the up-wave operation re-engages to expand the working set further. This cycle continues until the entire array is sorted. The up-wave thus serves as the main driver and entry point for the overall wave sort process. The mechanics of the up-wave process are illustrated in Fig. 1, and Listing 1.
1type seq []int
2
3func (s seq) upwave(start, end int) {
@@
21}
22
23func (s seq) WaveSort() {
24 if len(s) < 2 {
25 return
26 }
27 s.upwave(0, len(s)-1)
28}
Listing 1: Up-wave Algorithm
@@ 2.2.1. Partitioning
Initialize pointer i to one position before the first element, and pointer j to one position after the last element. Let p be the pivot. Increment i until the element at index i is greater than or equal to the pivot. If i == j, return i as the middle index m. Then decrement j until the element at index j is less than or equal to the pivot. If j == i, return i as the middle index m. Swap the elements at indices i and j. Repeat this process. The partitioning process is illustrated in Fig. 2 and Listing  2.
1// l is the index of the left bound of the unsorted part.
2// r is the index of the right bound of the unsorted part which
3is the firt element of the sorted part.
4// p is the index of the pivot.
5func (s seq) partition(l, r, p int) int {
@@
28}
Listing 2: Partitioning Algorithm
@@ 2.2.2. Block swapping
If the left block is shorter than the right block, define its length as $ll=r-m$, where r is the index of the first element in the right block, and m is the index of the first element in the left block. Set $nm=p-ll+1$, and initialize $j=m$. Store the element at index m in a temporary variable. Move the element at index $j+ll$ to index j, then update j to $j+ll$. If $j\geq nm$, replace the element at index $j+ll$ with the one at index $j-nm+m$. Repeat this process for $p-m+1$ iterations, as shown in Listing  3.
@@
Listing 3: Block-swap Algorithm sl
If the left block is not shorter than the right one, swap elements between the right and left blocks one by one. Only a portion of the unsorted block, equal in size to the sorted block, is moved, as shown in Listing  4.
1// m is the index of the first element of the left block.
2// r is the index of the first element of the right block.
3// p is the index of the last element of the right block.
4func (s seq) blockSwap(m, r, p int) {
@@
16}
17
18func (s seq) blockSwap_sr(m, r, p int) {
19 i := m
20 tmp := s[I]
21 j := r
22 for j < p {
23 s[i] = s[j]
24 i++
25 s[j] = s[i]
26 j++
27 }
28 s[i] = s[j]
29 s[j] = tmp
30}
Listing 4: Block-swap Algorithm
@@
The same partition and block swap operations are recursively applied to these new sections, progressively forming smaller subsections with the same structure. This continues until either the sorted or unsorted portion is fully exhausted. If the sorted portion is exhausted, an up-wave can be initiated to process the remaining unsorted portion, starting from its rightmost elements. If the unsorted portion is exhausted, the sorting process is complete. The down-wave procedure is illustrated in Fig. 4 and Listing  5.
@@
Listing 5: Down-wave
@@ 4. Experimental Results
Experimental code was run on randomly shuffled sequences and compared to quicksort. Results demonstrate that the basic wave sort algorithm performs close to the theoretical minimum comparison limit of $log\_2n!$. Across multiple experiments, wave sort required approximately $1\%$ more comparisons than this theoretical limit on average. This efficiency gap decreases as sequence length increases.
@@
Though wave sort requires fewer comparisons than quicksort, it performs more data swaps and requires additional block swaps. In wave sort, each block swap moves each element once and minimizes access to the temporary variable to only once in most cases. Each swap is equivalent to approximately 3 block swaps. The total data movement in basic wave sort is more than twice that of quicksort.
@@
In practice, many sequences encountered are presorted or partially presorted. Wave sort exhibits its worst-case performance when processing fully presorted sequences.
@@ 5. Adaptive Improvement
For sorted or partially sorted sequences, regardless of whether they are reversed, an additional check is incorporated into the up-wave to identify the longest presorted sequence as the sorted portion. For a fully sorted sequence, this results in only $n-1$ comparisons. If the sequence is completely reversed, the number of swaps is reduced to $\frac{n}{2}$. This modification effectively short-circuits the processing of presorted sequences (see Listing  6).
@@
Listing 6: Adaptive Wave Sort
@@
As the adaptive approach adds one more step to check the boundary of the sorted sequence, the number of comparisons increases slightly. This approach trades off swaps for block swaps. The comparisons increase by approximately $0.6\%$. The number of swaps decreases by approximately $1\%$, while block swaps decrease by approximately $7\%$.
@@ 6. Trade off Comparison for Swap
Since wave sort performs more than twice the swaps compared to quicksort, there is room for optimization if such trade-offs prove beneficial. The basic strategy is to increase the expansion rate in the up-wave so that partitioning works more effectively to reduce swaps. In both down-wave and up-wave processes, sorting short sequences with insertion sort can replace swaps with block swaps for greater efficiency. The smaller sorted portions also reduce the number of block swaps required. An implementation showing these trade-offs is present in Listing  7
@@
Listing 7: Trade-off Comparison
A typical trade-off result is shown in Table 6 with the implementation demonstrated in the code above. One comparison trades for approximately 4 swaps plus 80 block swaps, which is equivalent to approximately 31 swaps. More detailed mathematical analysis is required to determine the optimal expansion rate. Generally, the longer the sorted portion, the larger the expansion rate should be, based on the probability equation provided above.

### v2

@@ 2. Core Algorithm
2.1. Driver function
This function serves as the initial entry point, handling the base case for sequences with fewer than two elements by simply returning. For sequences of two or more elements, it delegates the core sorting logic to the core algorithm.
1type seq []int
2
3func (s seq) WaveSort() {
4 if len(s) < 2 {
5 return
6 }
7 s.upwave(0, len(s)-1)
8}
Listing 1: Driver function
@@ 2.1. Driver function
Once the down-wave completes sorting this set, thereby enlarging the overall sorted region, the up-wave operation re-engages to expand the working set further. This cycle continues until the entire array is sorted. The up-wave thus serves as the main driver and entry point for the overall wave sort process. The mechanics of the up-wave process are illustrated in Fig. 1, and Listing 2.
1func (s seq) upwave(start, end int) {
@@
19}
Listing 2: Up-wave Algorithm
@@ 2.3.1. Partitioning
Initialize pointer i to one position before the first element, and pointer j to one position after the last element. Let p be the pivot. Increment i until the element at index i is greater than or equal to the pivot. If i == j, return i as the middle index m. Then decrement j until the element at index j is less than or equal to the pivot. If j == i, return i as the middle index m. Swap the elements at indices i and j. Repeat this process. The partitioning process is illustrated in Fig. 2 and Listing  3.
1func (s seq) partition(l, r, p int) int {
@@
24}
Listing 3: Partitioning Algorithm
@@ 2.3.2. Block swapping
If the left block is shorter than the right block, define its length as $ll=r-m$, where r is the index of the first element in the right block, and m is the index of the first element in the left block. Set $nm=p-ll+1$, and initialize $j=m$. Store the element at index m in a temporary variable. Move the element at index $j+ll$ to index j, then update j to $j+ll$. If $j\geq nm$, replace the element at index $j+ll$ with the one at index $j-nm+m$. Repeat this process for $p-m+1$ iterations, as shown in Listing  4.
@@
Listing 4: Short left
If the left block is not shorter than the right one, swap elements between the right and left blocks one by one. Only a portion of the unsorted block, equal in size to the sorted block, is moved, as shown in Listing  6.
1func (s seq) blockSwap_sr(m, r, p int) {
2 i := m
3 tmp := s[I]
4 j := r
5 for j < p {
6 s[i] = s[j]
7 i++
8 s[j] = s[i]
9 j++
10 }
11 s[i] = s[j]
12 s[j] = tmp
13}
Listing 5: Short right
All together, the two scenarios are decided by the function shown in Listing  6.
1func (s seq) blockSwap(m, r, p int) {
@@
13}
Listing 6: blockSwap
@@
The same partition and block swap operations are recursively applied to these new sections, progressively forming smaller subsections with the same structure. This continues until either the sorted or unsorted portion is fully exhausted. If the sorted portion is exhausted, an up-wave can be initiated to process the remaining unsorted portion, starting from its rightmost elements. If the unsorted portion is exhausted, the sorting process is complete. The down-wave procedure is illustrated in Fig. 4 and Listing  7.
@@
Listing 7: Down-wave
@@ 4. Experimental Results
Experimental code was run on randomly shuffled sequences and compared to quicksort. Results demonstrate that the basic wave sort algorithm performs close to the theoretical minimum comparison limit of $log_{2}(n!)$. Across multiple experiments, wave sort required approximately $1\%$ more comparisons than this theoretical limit on average. This efficiency gap decreases as sequence length increases (See table 1).
@@
Though wave sort requires fewer comparisons than quicksort, it performs more data swaps and requires additional block swaps. In wave sort, each block swap moves each element once and minimizes access to the temporary variable to only once in most cases. Each swap is equivalent to approximately 3 block swaps. The total data movement in basic wave sort is more than twice that of quicksort (See table 2).
@@
In practice, many sequences encountered are presorted or partially presorted. Wave sort exhibits its worst-case performance when processing fully presorted sequences (See table 3).
@@ 5. Adaptive Improvement
For sorted or partially sorted sequences, regardless of whether they are reversed, an additional check is incorporated into the up-wave to identify the longest presorted sequence as the sorted portion. For a fully sorted sequence, this results in only $n-1$ comparisons. If the sequence is completely reversed, the number of swaps is reduced to $\frac{n}{2}$. This modification effectively short-circuits the processing of presorted sequences (see Listing  8 and table 4).
@@
Listing 8: Adaptive Wave Sort
@@
As the adaptive approach adds one more step to check the boundary of the sorted sequence, the number of comparisons increases slightly. This approach trades off swaps for block swaps. The comparisons increase by approximately $0.6\%$. The number of swaps decreases by approximately $1\%$, while block swaps decrease by approximately $7\%$ (See table 5).
@@ 6. Trade off Comparison for Swap
Since wave sort performs more than twice the swaps compared to quicksort, there is room for optimization if such trade-offs prove beneficial. The basic strategy is to increase the expansion rate in the up-wave so that partitioning works more effectively to reduce swaps. In both down-wave and up-wave processes, sorting short sequences with insertion sort can replace swaps with block swaps for greater efficiency. The smaller sorted portions also reduce the number of block swaps required. An implementation showing these trade-offs is present in Listing  9
@@
Listing 9: Trade-off Comparison
A typical trade-off result is shown in Table 6 with the implementation demonstrated in the code above. One comparison trades for approximately 4 swaps plus 80 block swaps, which is equivalent to approximately 31 swaps (See table 6). More detailed mathematical analysis is required to determine the optimal expansion rate. Generally, the longer the sorted portion, the larger the expansion rate should be, based on the probability equation provided above.